# Faster Convergence of Riemannian Stochastic Gradient Descent with Increasing Batch Size

### v3

@@ Appendix E Additional Numerical Results
This section presents the numerical results for the warm-up LR. We used a constant BS, an exponential growth BS equation 5, a polynomial growth BS equation 6, and a warm-up LR with an increasing part (exponential growth LR) and a decaying part (a constant LR equation 1, a diminishing LR equation 2, a cosine annealing LR equation 3, or a polynomial decay LR equation 4). We set $K^{\prime}=200$. When we used an exponential growth LR, a polynomial growth BS was not required to satisfy $\delta^{2l}<\gamma$, but an exponential growth BS was required to satisfy it (see Section 3.3). For comparison, even when using a polynomial growth BS, we used a setting that satisfies this condition. Furthermore, we chose $\eta_{\max}$ (initial value of decaying part) from $\{0.5,0.05,0.005\}$. When using an exponential BS, we set initial batch size $b_{0}\coloneq 3^{4}$ in Cases A, B and $b_{0}\coloneq 3^{3}$ in Cases C, D. When using the other BSs, we set the same initial batch size with Section 4. From the definition of exponential growth LR equation 7, we can represent the part after warm-up as $\eta_{T_{w}-1}=\eta_{0}\delta^{l_{w}}$. Consequently, we chose hyperparameters $(l,l_{w},\gamma,\eta_{\max},\eta_{0})$ satisfying $\eta_{0}\delta^{l_{w}}=\eta_{\max}$ and $\delta^{2l}<\gamma$; i.e., $\delta=(\frac{\eta_{\max}}{\eta_{0}})^{\frac{1}{l_{w}}}<\gamma^{\frac{1}{2l}}$. Hence, when $l=l_{w}=3$ and $\gamma=3.0$, we set $\eta_{\max}=0.5,0.05,0.005$ and $\eta_{0}=\frac{5}{17},\frac{5}{170},\frac{5}{1700}$ respectively. In this setting, $\delta=\sqrt[3]{\frac{\eta_{\max}}{\eta_{0}}}=\sqrt[3]{1.7}<\sqrt[6]{3}=\sqrt[6]{\gamma}=\gamma^{\frac{1}{2l}}$ holds. And, when $l=3,l_{w}=8,\gamma=3$, we set $\eta_{\max}=0.5,0.05,0.005$, $\eta_{0}=\frac{1}{8},\frac{1}{80},\frac{1}{800}$ respectively. In this setting, $\delta=\sqrt[8]{\frac{\eta_{\max}}{\eta_{0}}}=\sqrt[8]{4}<\sqrt[6]{3}=\sqrt[6]{\gamma}=\gamma^{\frac{1}{2l}}$ holds. Since we terminated RSGD after the $3000$th iteration and set $K^{\prime}=200$, $l=l_{w}=3$ (resp. $l=3,l_{w}=8$). Those settings mean 5 times BS increasing and 3 times LR increasing (resp. 5 times BS increasing and 8 times LR increasing).
E.1 Principle Components Analysis
@@
Figures 11 and 13 plot performance in terms of the gradient norm of the objective function against the number of iterations for a warm-up LR with decay parts equation 1, equation 2, equation 3, and equation 4 on the COIL100 and MNIST datasets, respectively. Figures 12 and 14 plot performance in terms of the objective function value against the number of iterations for a warm-up LR with decay parts equation 1, equation 2, equation 3, and equation 4 for the COIL100 and MNIST datasets, respectively.
Figure 11: Norm of the gradient of the objective function against number of iterations for warm-up LRs that have an increasing part with three increments on COIL100 dataset.
Figure 12: Objective function value against number of iterations for warm-up LRs that have an increasing part with three increments on COIL100 dataset.
Figure 13: Norm of the gradient of the objective function against number of iterations for warm-up LRs that have an increasing part with three increments on MNIST dataset.
Figure 14: Objective function value against number of iterations for warm-up LRs that have an increasing part with three increments on MNIST dataset.
@@
Figures 15 and 17 plot the performance in terms of the gradient norm of the objective function against the number of iterations for a warm-up LR with decay parts equation 1, equation 2, equation 3, and equation 4 for the COIL100 and MNIST datasets, respectively. Figures 16 and 18 plot the performance in terms of the objective function value against the number of iterations of a warm-up LR with decay parts equation 1, equation 2, equation 3, and equation 4 for the COIL100 and MNIST datasets, respectively.
Figure 15: Norm of the gradient of the objective function against number of iterations for warm-up LRs that have an increasing part with eight increments on COIL100 dataset.
Figure 16: Objective function value against number of iterations for warm-up LRs that have an increasing part with eight increments on COIL100 dataset.
Figure 17: Norm of the gradient of the objective function against number of iterations for warm-up LRs that have an increasing part with eight increments on MNIST dataset.
Figure 18: Objective function value against number of iterations for warm-up LRs that have an increasing part with eight increments on MNIST dataset.
E.2 Low-rank Matrix Completion
@@
Figures 19 and 21 plot the performance in terms of the gradient norm of the objective function against the number of iterations for a warm-up LR with decay parts equation 1, equation 2, equation 3, and equation 4 on the MovieLens-1M and Jester datasets, respectively. Figures 19 and 22 plot the performance in terms of the objective function value against the number of iterations for a warm-up LR with decay parts equation 1, equation 2, equation 3, and equation 4 on the MovieLens-1M and Jester datasets, respectively.
Figure 19: Norm of the gradient of the objective function against number of iterations for warm-up LRs that have an increasing part with three increments on MovieLens-1M dataset.
Figure 20: Objective function value against number of iterations for warm-up LRs that have an increasing part with three increments on MovieLens-1M dataset.
Figure 21: Norm of the gradient of the objective function against number of iterations for warm-up LRs that have an increasing part with three increments on Jester dataset.
Figure 22: Objective function value against number of iterations for warm-up LRs that have an increasing part with three increments on Jester dataset.
@@
Figures 23 and 25 plot the performance in terms of the gradient norm of the objective function against the number of iterations for a warm-up LR with decay parts equation 1, equation 2, equation 3, and equation 4 on the MovieLens-1M and Jester datasets, respectively. Figures 23 and 26 plot the performance in terms of the objective function value against the number of iterations for a warm-up LR with decay parts equation 1, equation 2, equation 3, and equation 4 on the MovieLens-1M and Jester datasets, respectively.
Figure 23: Norm of the gradient of the objective function against number of iterations for warm-up LRs that have an increasing part with eight increments on MovieLens-1M dataset.
Figure 24: Objective function value against number of iterations for warm-up LRs that have an increasing part with eight increments on MovieLens-1M dataset.
Figure 25: Norm of the gradient of the objective function against number of iterations for warm-up LRs that have an increasing part with eight increments on Jester dataset.
Figure 26: Objective function value against number of iterations for warm-up LRs that have an increasing part with eight increments on Jester dataset.

### v4

@@ Appendix G Additional Numerical Results
This section presents the numerical results for a warm-up LR. We used a constant BS, an exponential growth BS (5), a polynomial growth BS (6), and a warm-up LR with an increasing part (exponential growth LR) and a decaying part (either a constant LR (1), a diminishing LR (2), a cosine annealing LR (3), or a polynomial decay LR (4)). We set $K^{\prime}=200$. When we used an exponential growth LR, a polynomial growth BS was not required to satisfy $\delta^{2l}<\gamma$, whereas an exponential growth BS was required to satisfy it (see Section 3.3). For comparison, even when using a polynomial growth BS, we adopted a setting that satisfies this condition. Furthermore, we chose $\eta_{\max}$ (the initial value of decaying part) from $\{0.5,0.05,0.005\}$. When using an exponential BS, we set the initial BS $b_{0}\coloneq 3^{4}$ in Cases A, B and $b_{0}\coloneq 3^{3}$ in Cases C, D. When using the other BSs, we set the same initial batch size as in Section 4.
From the definition of an exponential growth LR (7), we can represent the part after warm-up as $\eta_{T_{w}-1}=\eta_{0}\delta^{l_{w}}$. Consequently, we chose hyperparameters $(l,l_{w},\gamma,\eta_{\max},\eta_{0})$ satisfying $\eta_{0}\delta^{l_{w}}=\eta_{\max}$ and $\delta^{2l}<\gamma$; i.e., $\delta=(\frac{\eta_{\max}}{\eta_{0}})^{\frac{1}{l_{w}}}<\gamma^{\frac{1}{2l}}$. Hence, when $l=l_{w}=3$ and $\gamma=3.0$, we set $\eta_{\max}=0.5,0.05,0.005$ and $\eta_{0}=\frac{5}{17},\frac{5}{170},\frac{5}{1700}$, respectively. In this setting, $\delta=\sqrt[3]{\frac{\eta_{\max}}{\eta_{0}}}=\sqrt[3]{1.7}<\sqrt[6]{3}=\sqrt[6]{\gamma}=\gamma^{\frac{1}{2l}}$ holds. Similarly, when $l=3,l_{w}=8,\gamma=3$, we set $\eta_{\max}=0.5,0.05,0.005$, $\eta_{0}=\frac{1}{8},\frac{1}{80},\frac{1}{800}$. respectively. In this setting, $\delta=\sqrt[8]{\frac{\eta_{\max}}{\eta_{0}}}=\sqrt[8]{4}<\sqrt[6]{3}=\sqrt[6]{\gamma}=\gamma^{\frac{1}{2l}}$ holds. Because we terminated RSGD after the $3000$th iteration and set $K^{\prime}=200$, we used $l=l_{w}=3$ (resp. $l=3,l_{w}=8$). In the first setting, the batch size increases five times and the learning rate three times; in the second, the batch size increases five times and the learning rate eight times.
G.1 Principle Components Analysis
@@
Figures 13 and 15 plot performance in terms of the gradient norm of the objective function versus the number of iterations for a warm-up LR with decay parts given by (1), (2), (3), and (4) on the COIL100 and MNIST datasets, respectively. Figures 14 and 16 plot performance in terms of the objective function value versus the number of iterations for a warm-up LR with decay parts given by (1), (2), (3), and (4) for the COIL100 and MNIST datasets, respectively.
Figure 13: Norm of the gradient of the objective function versus number of iterations for warm-up LRs that have an increasing part with three increments on COIL100 dataset (PCA).
Figure 14: Objective function value (loss) versus number of iterations for warm-up LRs that have an increasing part with three increments on COIL100 dataset (PCA).
Figure 15: Norm of the gradient of the objective function versus number of iterations for warm-up LRs that have an increasing part with three increments on MNIST dataset (PCA).
Figure 16: Objective function value (loss) versus number of iterations for warm-up LRs that have an increasing part with three increments on MNIST dataset (PCA).
@@
Figures 17 and 19 plot performance in terms of the gradient norm of the objective function versus the number of iterations for a warm-up LR with decay parts given by (1), (2), (3), and (4) for the COIL100 and MNIST datasets, respectively. Figures 18 and 20 plot performance in terms of the objective function value versus the number of iterations for a warm-up LR with decay parts given by (1), (2), (3), and (4) for the COIL100 and MNIST datasets, respectively.
Figure 17: Norm of the gradient of the objective function versus number of iterations for warm-up LRs that have an increasing part with eight increments on COIL100 dataset (PCA).
Figure 18: Objective function value (loss) versus number of iterations for warm-up LRs that have an increasing part with eight increments on COIL100 dataset (PCA).
Figure 19: Norm of the gradient of the objective function versus number of iterations for warm-up LRs that have an increasing part with eight increments on MNIST dataset (PCA).
Figure 20: Objective function value (loss) versus number of iterations for warm-up LRs that have an increasing part with eight increments on MNIST dataset (PCA).
G.2 Low-rank Matrix Completion
@@
Figures 21 and 23 plot performance in terms of the gradient norm of the objective function versus the number of iterations for a warm-up LR with decay parts given by (1), (2), (3), and (4) on the MovieLens-1M and Jester datasets, respectively. Figures 21 and 24 plot performance in terms of the objective function value versus the number of iterations for a warm-up LR with decay parts given by (1), (2), (3), and (4) on the MovieLens-1M and Jester datasets, respectively.
Figure 21: Norm of the gradient of the objective function versus number of iterations for warm-up LRs that have an increasing part with three increments on MovieLens-1M dataset (LRMC).
Figure 22: Objective function value (loss) versus number of iterations for warm-up LRs that have an increasing part with three increments on MovieLens-1M dataset (LRMC).
Figure 23: Norm of the gradient of the objective function versus number of iterations for warm-up LRs that have an increasing part with three increments on Jester dataset (LRMC).
Figure 24: Objective function value (loss) versus number of iterations for warm-up LRs that have an increasing part with three increments on Jester dataset (LRMC).
@@
Figures 25 and 27 plot performance in terms of the gradient norm of the objective function versus the number of iterations for a warm-up LR with decay parts given by (1), (2), (3), and (4) on the MovieLens-1M and Jester datasets, respectively. Figures 25 and 28 plot performance in terms of the objective function value versus the number of iterations for a warm-up LR with decay parts given by (1), (2), (3), and (4) on the MovieLens-1M and Jester datasets, respectively.
Figure 25: Norm of the gradient of the objective function versus number of iterations for warm-up LRs that have an increasing part with eight increments on MovieLens-1M dataset (LRMC).
Figure 26: Objective function value (loss) versus number of iterations for warm-up LRs that have an increasing part with eight increments on MovieLens-1M dataset (LRMC).
Figure 27: Norm of the gradient of the objective function versus number of iterations for warm-up LRs that have an increasing part with eight increments on Jester dataset (LRMC).
Figure 28: Objective function value (loss) versus number of iterations for warm-up LRs that have an increasing part with eight increments on Jester dataset (LRMC).